# Boundary layers in thermal convection are fluctuation-dominated

### v1

@@ VI Conclusions
Our DNS of the turbulent Rayleigh-Bénard convection encompasses a Cartesian domain with $\Gamma=4$ and periodic boundary conditions in the horizontal directions. These simulations up to $Ra=10^{11}$ are aimed at being close to the original canonical case of a plane convection layer between a pair of infinitely extended rigid plates. We demonstrated that a standard mean flow analysis with a subsequent match to laminar boundary layer profiles gives no systematic profiles; to the extent that we can define them, they are very small in shear Reynolds numbers. In the long-time limit, which we could follow for 1600 $T_{f}$ at $Ra=10^{9}$, the velocity mean profiles have to converge to $\mbox{lim}_{t\to 0}\langle u_{i}\rangle_{A,t}(z)\to 0$ due to homogeneity in $x$ and $y$. The simulations by Hartlep et al. (2003) (for $Ra\leq 1\times 10^{7}$) also found that the mean flow contained very little energy, but De et al. (2018) found a long time periodicity to the mean flow for $Ra\leq 2\times 10^{6}$.
Rather than having a mean flow profile with small fluctuations, we are faced with small mean velocity amplitudes with fluctuations that are up to 3 orders of magnitude larger when the statistics are taken over finite time intervals $\Delta t$, as seen from comparisons of table 2 with the data in figure 4. The strong fluctuations cause the boundary layer thickness scales of temperature and velocity to differ by an order of magnitude for the highest Rayleigh numbers, as summarized in table 3. These results hold for all Rayleigh numbers of the series, but become particularly pronounced for $Ra\geq 10^{9}$, a range to which most previous larger-aspect-ratio DNS studies did not advance.
Our analysis also revealed that the velocity boundary layer in the present configuration is a carpet of differently oriented shear dominated patches interspersed by regions of incoherent mean flow, the latter of which occupies about 60% of the plate area for all Rayleigh numbers. This heterogeneous composition crystallizes particularly for $Ra\gtrsim 10^{9}$, underlying again the importance of DNS with larger aspect ratios and high Rayleigh numbers. The present results also raise many questions on the possible transition mechanisms of the boundary layer to a turbulent regime and the possible consequences for the global heat transfer. The heterogeneity of the velocity boundary layer detected here suggests to us the prevalence of local, rather than global, instability mechanisms, which would bring us back to the marginal stability concept of the boundary layer, see e.g. Howard (1966). However, figure 2(c) shows a power law fit of $Ra_{\delta,{\rm rms}}=ARa^{\gamma}$ with a very small exponent $\gamma=0.077$ and $A\approx 41$. The resulting $Ra_{\delta,{\rm rms}}$ are by at least a factor of 4 smaller than Howard’s critical Rayleigh number of $Ra_{\delta}\sim{\cal O}(10^{3})$. These open points will form the subject of further studies including higher Rayleigh numbers in this configuration.
Acknowledgements.
@@
The work of RJS is funded by the European Union (ERC, MesoComp, 101052786). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. The work of JDS was supported by a Mercator Fellowship of the Deutsche Forschungsgemeinschaft within the Priority Programme DFG-SPP 1881 on Turbulent Superstructures. The authors gratefully acknowledge the Gauss Centre for Supercomputing e.V. (https://www.gauss-centre.eu) for funding this project by providing computing time through the John von Neumann Institute for Computing (NIC) on the GCS Supercomputer JUWELS at Jülich Supercomputing Centre (JSC).

### v2

@@ 7 Final discussion
Our DNS of the turbulent Rayleigh-Bénard convection encompasses a Cartesian domain with $\Gamma=4$, with no-slip horizontal walls and periodic boundary conditions for the side faces. These simulations up to $Ra=10^{11}$ are aimed to approach the original canonical case of a plane convection layer between a pair of infinitely extended rigid plates. We demonstrated that a standard mean flow profile (obtained by combining averages with respect to time and the entire horizontal cross sectional plane) have very small magnitudes, and that efforts to match them to laminar boundary layer profiles produced no conclusive results. To the extent that we can define the boundary layers, they give very small shear Reynolds numbers (see below). In the long-time limit, which we have followed for 1600 $T_{f}$ at $Ra=10^{9}$, the velocity mean profiles have to converge to $\mbox{lim}_{t\to\infty}\langle u_{i}\rangle_{A,t}(z)\to 0$ due to statistical homogeneity in $x$ and $y$. The simulations by Hartlep et al. (2003) (for $Ra\leq 1\times 10^{7}$) also showed that the mean flow contained very little kinetic energy, but De et al. (2018) found a long time periodicity in the mean flow for low Rayleigh numbers, $Ra\leq 2\times 10^{6}$.
Rather than having a mean flow profile with small velocity fluctuations, we are faced with small mean velocity amplitudes in the presence of velocity fluctuations that are up to 2 orders of magnitude larger when the statistics are taken over finite time intervals $\tau_{\rm total}$, as seen from comparisons of table 2 with the data in figure 5. This central result also holds when the aspect ratio of the simulation is varied. It is our view that fluctuations will be relevant for all configurations which includes closed cells of $\Gamma\lesssim 1$, see e.g. figure 9. But their relevance is strongest in the statistically homogeneous plane layer with periodic boundary conditions in the horizontal direction—the configuration that comes closest to the original physical problem of turbulent convection (Spiegel, 1962), as relevant for most geo- and astrophysical applications.
We also showed that the corresponding shear Reynolds numbers, which are based on mean flow quantities for a finite averaging time, remain very small because the characteristic velocities $U_{\infty}$ are small. The strong fluctuations cause the fluctuation thicknesses of temperature and velocity (defined as the near-wall maxima of the rms profiles) to differ by an order of magnitude for the highest Rayleigh numbers, as summarized in table 3. This difference increases with Rayleigh number (though the Prandtl number is held fixed at order unity); it becomes particularly pronounced for $Ra\geq 10^{9}$, a range beyond which previous larger-aspect-ratio DNS studies rarely advanced.
Furthermore, our analysis revealed that the velocity boundary region in the present configuration is a carpet of differently oriented time-dependent and shear-dominated (coherent) regions interspersed by regions of incoherent flow. The latter regions occupy about 60% of the plate area for all Rayleigh numbers. This heterogeneous composition crystallizes particularly for $Ra\gtrsim 10^{9}$, underlying again the importance of DNS with larger aspect ratios and high Rayleigh numbers. The incoherent regions in the present flow can be as large as the entire cross section of a cylindrical cell at $\Gamma\sim 1$. The coherent regions are the near-wall footprint of the circulation rolls which form the large-scale turbulent superstructure pattern (Stevens et al., 2018; Pandey et al., 2018). They change their orientation continually and thus result in a net zero mean flow as stated above.
Finally, we showed that the velocity fluctuation thicknesses decrease when they are conditioned on shear-dominated patches in the near-wall region. They are then closer to those scales which are obtained in turbulent convection in closed cylindrical cells of aspect ratio $\Gamma\lesssim 1$. The geometry of the closed cell enforces a large-scale circulation which is mostly shear-dominated in the vicinity of the walls, as already shown in Schumacher et al. (2016) where the time-dependence of the orientation has been eliminated. This causes smaller velocity thickness scales that are, however, still larger than the thermal boundary layer thickness at $Pr\sim 1$. Nevertheless, the fluctuation thickness is the consistently definable velocity boundary layer scale for the present turbulent convection flow.
The present results also raise many questions on the possible transition mechanisms of the boundary layer to a turbulent regime and the possible consequences for the global heat transfer. Differently from wall-bounded shear flows, we detect velocity fluctuations everywhere, even though at different strengths; see again figure 9. Furthermore, we do not observe a Rayleigh number dependence of the ratio of coherent (“laminar”) to incoherent (“turbulent”) regions. The time scales, at which these complex spatio-temporal patterns change, become increasingly shorter with increasing Rayleigh number. A variation of the threshold for the decomposition into coherent and incoherent boundary regions practically did not affect this Rayleigh number independence.
The spatio-temporal heterogeneity of the velocity boundary layer, which we detected here, suggests to us the prevalence of local, rather than global, instability mechanisms, which would bring us back to the marginal stability concept of the boundary layer, see e.g. Howard (1966) and for a detailed boundary layer model with plume formation, Theerthan & Arakeri (1998). However, figure 10 shows a power law fit of $Ra_{\delta,{\rm rms}}=ARa^{\gamma}$ with a very small exponent $\gamma=0.077$ and $A\approx 41$. The resulting $Ra_{\delta,{\rm rms}}$ are by at least a factor of 4 smaller than Howard’s critical Rayleigh number of $Ra_{\delta}\sim{\cal O}(10^{3})$. Additional data from experiments show that even at $Ra\sim 10^{17}$ a thickness-based Rayleigh number $Ra_{\delta}$ barely reaches a value of $10^{3}$. This challenges the original marginal stability concept. This question and higher Rayleigh number simulation in the present configuration form the subject of further study.
Acknowledgements
The authors thank Jaywant Arakeri, Björn Hof, Hassan Nagib and Peter Schmid for helpful discussions.
Funding
@@ Acknowledgements
The work of RJS is funded by the European Union (ERC, MesoComp, 101052786). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. The work of JDS was supported by a Mercator Fellowship of the Deutsche Forschungsgemeinschaft within the Priority Programme DFG-SPP 1881 on Turbulent Superstructures. The authors gratefully acknowledge the Gauss Centre for Supercomputing e.V. (https://www.gauss-centre.eu) for funding the project nonbou by providing computing time through the John von Neumann Institute for Computing (NIC) on the GCS Supercomputer JUWELS Booster at Jülich Supercomputing Centre (JSC).
Declaration of interests
The authors report no conflict of interest.
Data availability statement
The data that support the findings of this study are available on reasonable request.
Author ORCIDs
Roshan J. Samuel https://orcid.org/$0$-0002-1280-9881
Mathis Bode https://orcid.org/$0$-0001-9922-9742
Janet D. Scheel https://orcid.org/$0$-0002-1669-4188
Katepalli R. Sreenivasan https://orcid.org/$0$-0002-3943-6827
Jörg Schumacher https://orcid.org/$0$-0002-1359-4536
Author contributions
All authors designed the research. R.J.S., M.B. and J.S. carried out the supercomputer simulations. R.J.S., J.D.S and J.S. analysed the simulation data and generated the figures. All authors wrote the manuscript.